# MMedFD: A Real-world Healthcare Benchmark for Multi-turn Full-Duplex Automatic Speech Recognition

### v1

@@ 4 Results
Benchmark Multi- Turn Full- Duplex Health- Care Nature Roles Dur. (h) #Dialogs #Turns Language Evaluation MultiMed[3] $\times$ $\times$ Real-world 6 150 Multiling. ASR(WER/CER) VietMed[17] $\times$ $\times$ Real-world 6 16 Vietnamese ASR(WER) PriMock57[6] $\times$ Simulated 2 9 57 $\approx$5244 English ASR(WER), Task Fareez et al.[18] $\times$ Simulated 2 55.15 272 English ASR(WER), Human-Eval myMediCon[19] $\times$ $\times$ Read 2 11 Burmese ASR(WER) AfriSpeech-200[20] $\times$ $\times$ Mixed 1 $\approx$123 Afr. English ASR(WER) SpokenWOZ[21] $\times$ $\times$ Real-world 2 249 5,700 203,000 English ASR(WER), Task VoxDialogue[22] $\times$ $\times$ Mixed 2 42.56 4,500 30,700 English LLM-Eval MTalk-Bench[7] $\times$ $\times$ Mixed 2 2.45 90 568 English S2S, LLM-Eval, Human-Eval MMedFD (ours) Real-world 2 136.91 1,805 10,814 Chinese ASR(WER/MC-WER), LLM-Eval Notes: The symbols ✓ and $\times$ denote presence and absence, respectively, “–” indicates information not reported. Definitions: Multi-turn indicates explicit turn structure with cross-turn context, Full-duplex refers to user–agent settings in which the system speaks while continuously listening, Healthcare marks corpora collected in clinical/medical contexts. Nature categorizes collection modality as Real-world (naturally occurring audio), Simulated (scripted/acted clinical dialogues), Read (read speech), or Mixed (combined sources). Scale: Dur. (h) denotes total hours, #Dialogs counts conversational sessions, #Turns counts utterance-level turns, Roles indicates the number of distinct speaker roles. Evaluation abbreviations: ASR (Automatic Speech Recognition), WER (Word Error Rate), CER (Character Error Rate), MC-WER (medical-content WER), Task (objective task metrics such as dialogue-state accuracy, slot F1, task success), LLM-Eval (Large Language Model–based evaluation), Human-Eval (human evaluation), TTS (Text-to-Speech evaluation), and S2S (Speech-to-Speech evaluation via arena/rubrics).

### v2

@@ 4 Results
Benchmark Multi- Turn Full- Duplex Health- Care Nature Roles Dur. (h) #Dialogs #Turns Language Evaluation MultiMed[3] $\times$ $\times$ Real-world 6 150 Multiling. ASR(WER/CER) VietMed[17] $\times$ $\times$ Real-world 6 16 Vietnamese ASR(WER) PriMock57[6] $\times$ Simulated 2 9 57 $\approx$5244 English ASR(WER), Task Fareez et al.[18] $\times$ Simulated 2 55.15 272 English ASR(WER), Human-Eval myMediCon[19] $\times$ $\times$ Read 2 11 Burmese ASR(WER) AfriSpeech-200[20] $\times$ $\times$ Mixed 1 $\approx$123 Afr. English ASR(WER) SpokenWOZ[21] $\times$ $\times$ Real-world 2 249 5,700 203,000 English ASR(WER), Task VoxDialogue[22] $\times$ $\times$ Mixed 2 42.56 4,500 30,700 English LLM-Eval MTalk-Bench[7] $\times$ $\times$ Mixed 2 2.45 90 568 English S2S, LLM-Eval, Human-Eval MMedFD (ours) Real-world 2 136.91 1,805 10,814 Chinese ASR(WER/HC-WER), LLM-Eval Notes: The symbols ✓ and $\times$ denote presence and absence, respectively, “–” indicates information not reported. Definitions: Multi-turn indicates explicit turn structure with cross-turn context, Full-duplex refers to user–agent settings in which the system speaks while continuously listening, Healthcare marks corpora collected in clinical/medical contexts. Nature categorizes collection modality as Real-world (naturally occurring audio), Simulated (scripted/acted clinical dialogues), Read (read speech), or Mixed (combined sources). Scale: Dur. (h) denotes total hours, #Dialogs counts conversational sessions, #Turns counts utterance-level turns, Roles indicates the number of distinct speaker roles. Evaluation abbreviations: ASR (Automatic Speech Recognition), WER (Word Error Rate), CER (Character Error Rate), HC-WER (Healthcare-content WER), Task (objective task metrics such as dialogue-state accuracy, slot F1, task success), LLM-Eval (Large Language Model–based evaluation), Human-Eval (human evaluation), TTS (Text-to-Speech evaluation), and S2S (Speech-to-Speech evaluation via arena/rubrics).